OpenStudy (anonymous):

Worlds population in 1960 was approximately 3 billion; in 2000 it was approximately 6 billion. Model this using a differential equation of your choice. You should include your reasons for your choice of models, the solution to the equation, and a graph spanning the given data and some time into the future. Please help me get this finished!!! Medal when done! :)

OpenStudy (ghazi):

i am sorry i could have solved but its a lengthy process , all i want to say is ..total stochastic calculus failed in developing and framing expansion of population so i am sure this equation wont give solution either

OpenStudy (unklerhaukus):

i think more than two data point would be helpful

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

@UnkleRhaukus Two points is enough it you assume it is some exponential function.

OpenStudy (unklerhaukus):

which ones can we assume?

OpenStudy (anonymous):

\[ P(t) = P_0e^{kt} \]\[ P(1960) = 3 = P_0e^{k(1960)}\\ P(2000) = 6 = P_0e^{k(2000)} \]Two equations, two variables \(P_0\) and \(k\).

OpenStudy (anonymous):

You could also think of it as, the double life of the function is 40 years, and suppose the function starts at year 1960. In that case: \[ P(t) = 3\cdot 2^{t/40} \]

OpenStudy (unklerhaukus):

but these models are so inaccurate

OpenStudy (anonymous):

What do you think the the carrying capacity as a function of time?

OpenStudy (anonymous):

Okay how can I make them accurate then?

OpenStudy (anonymous):

Because Im lost...One person says this and the other says something else...

OpenStudy (zehanz):

Accuracy is not the point here. When so little data is given, the only thing you can reasonably do is assume there is exponential growth. The data support @wio's formula: \(P(t)=3 \cdot 2^{\frac{1}{40}t}=3 \cdot e^{\ln2^{\frac{1}{40}t}}=3e^{\frac{\ln 2}{40} \cdot t} \approx 3e^{0.01733t}\). If we take the derivative of P, we get: \(P'(t)=0.01733P(t)\). The differential equation that fits the situation is therefore: \(\dfrac{dx}{dt}=0.01733x\). Solving can be done with separation of variables: \(\dfrac{dx}{x}=0.01733dt\) \(\int \dfrac{dx}{x}=\int 0.01733dt +C\); C is a real constant. \(ln|x|=0.01733t+C\) \(|x|=e^{0.01733t+C}=e^{C}e^{0.01733t}=C_{1}e^{0.01733t};~C_{1}\) is a positive constant. Because x is the population, it is always positive, so we don't need the absolute value: \(x=C_{1}e^{0.01733t}\). To find the constant, we set x(0)=3, so the constant is 3. Solution: \(x(t)=3e^{0.01733t}\), of course the same as the population function P. BTW: if we forget about the practical situation this is about, the general solution of the differential equation is \(x(t)=Ce^{0.01733t}\), where C can be any real constant. Here is also a graph with prediction for 2020:

OpenStudy (anonymous):

Okay thanks so much! So may I ask...What would be different if you would have made the choice of a different model?

OpenStudy (anonymous):

Please help me with this integration!

OpenStudy (anonymous):

If you wanted to change the model, you'd probably want the change of population (derivative of population) to tend toward \(0\) as the the population reached some carrying capacity.